A lot consists of (6+a) articles, 3 with minor defects, and 3 with major defects. The ‘a’

articles are chosen at random from the lot without replacement. Compute the

probability by using basic probability concepts, both articles are good.


1
Expert's answer
2021-08-10T11:55:16-0400

A lot consists of (6+a) articles, 6 with defects.

"a" articles are chosen at random from the lot without replacement.

Suppose that "X" be the number of articles with defects from "a" articles chosen


"P(X=0)=\dfrac{\dbinom{a}{a}\dbinom{6}{0}}{\dbinom{6+a}{a}}=\dfrac{6!a!}{(6+a)!}"

If "2" articles are chosen at random from the lot without replacement "(a\geq2)."

Suppose that "Y" be the number of articles with defects from "2" articles chosen


"P(Y=0)=\dfrac{\dbinom{a}{2}\dbinom{6}{0}}{\dbinom{6+a}{2}}=\dfrac{a!6!a!}{2!(a-2)!(6+a)!}"